Question 32 pts

Continuing with the example from Questions 1 and 2, you determine that the value of out$sigma is 3.0. Using the formula for relative efficiency given in class, you estimate that the MSE for the same trial conducted as a CRD would have been 13.5. The Relative Efficiency of the RBD compared to the CRD is therefore:

Group of answer choices

1.5

4.5

1.225

None of the above
Answer #1
relative efficiency of Θ̂ 2 to Θ̂ 1 is defined as = MSE(Θ̂ 1)/MSE(Θ̂ 2)

therefore  The Relative Efficiency of the RBD compared to the CRD is =(13.5/3^2) =1.5